# Stress out of charmonia

### v1

@@ IV.2 Gravitational form factor $D$: kinetic part
The GFF $D(q^{2}_{\perp})$ and the mass squared form factor $M^{2}(q^{2}_{\perp})$ can be extracted from either the light-front energy density $T^{+-}$, or the transverse stress $T^{11}+T^{22}$ and $T^{12}$. From the covariant light-front dynamics analysis [49], $T^{11}+T^{22}$ and $T^{12}$ are contaminated by spurious form factors and may violate the von Laue condition [^1] We recently discovered a new symmetry that may protect $T^{12}$ from the contamination of the spurious form factor [167]. It is not clear whether this symmetry is applicable to the present system.. $T^{+-}$ is the light-front energy density and it contains the interaction. While this operator is more involved, it is constrained by energy conservation which ensures the von Laue condition even after model truncation. We therefore use $T^{+-}$ to extract GFF $D$. Our method is general enough and can be applied to other bound-state systems. The HME of $T^{+-}$ consists of two parts:
@@
In the Breit frame $P_{\perp}=0$, only the invariant mass squared form factor $M^{2}(q^{2}_{\perp})$ survives. GFF $D$ is related to $M^{2}(q^{2}_{\perp})$ as,
$q_{\perp}^{2}D(q_{\perp}^{2})=2M^{2}(q^{2}_{\perp})-2(M^{2}+\frac{1}{4}q^{2}_{\perp})A(q^{2}_{\perp})\,.$ (65)
In Ref. [49], we propose to split $T^{+-}$ into two parts $T^{+-}=T^{+-}_{0}+T^{+-}_{\text{int}}$: a free part $T^{+-}_{0}$ and an interaction part $T^{+-}_{\text{int}}$. The corresponding mass squared form factor is also split into two pieces, $M^{2}(q^{2}_{\perp})=M^{2}_{0}(q^{2}_{\perp})+M^{2}_{\text{int}}(q^{2}_{\perp})$. The free energy density $T^{+-}_{0}$ is diagonal in Fock space, and the corresponding mass squared form factor admits an exact LFWF representation,
@@
The integrand is the $n$-body kinetic energy with an extra term $-\frac{1}{4}q^{2}_{\perp}$. This term accounts for the recoil effect, which has the same origin as the $D$-term as one can see from the free boson case [49].
@@
$M_{0}^{2}(q^{2}_{\perp})=\sum_{n}\sum_{\{s_{i}\}}\int\big{[}{\mathrm{d}}x_{i}{\mathrm{d}}^{2}r_{i\perp}\big{]}_{n}\widetilde{\psi}_{n}^{*}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\\ \times\sum_{j}e^{i\vec{r}_{j\perp}\cdot\vec{q}_{\perp}}\frac{-\nabla_{j\perp}^{2}+m_{j}^{2}-\frac{1}{4}q_{\perp}^{2}}{x_{j}}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\,.$ (67)
Its Fourier transform gives the OBD of the free mass squared,
$\mathcal{M}_{0}^{2}(r_{\perp})=\sum_{n}\sum_{\{s_{i}\}}\int\big{[}{\mathrm{d}}x_{i}{\mathrm{d}}^{2}r_{i\perp}\big{]}_{n}\sum_{j}\delta^{2}(\vec{r}_{j\perp}-\vec{r}_{\perp})\\ \times\bigg{\{}\widetilde{\psi}_{n}^{*}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\frac{-\nabla_{j\perp}^{2}+m_{j}^{2}}{x_{j}}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\\ +\frac{1}{4x_{j}}\nabla^{2}_{j\perp}\big{|}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\big{|}^{2}\bigg{\}}\,.$ (68)
@@ IV.3 Gravitational form factor $D$: interaction part
The interaction part of the EMT in coordinate space can be formally represented in Fock space using LFWFs. However, the expression is not particularly useful in quantum field theories, since it involves particle number changing. Furthermore, the interaction $v_{nm}$ also has to be renormalized. The effective interactions, such as the one-gluon-exchange we employed, are non-local in general. It may seem that a local interaction density has to be obtained from an underlying field theory. On the other hand, everything gravitates. As such, the EMT should exist for any system. As the dynamics is fully determined by the Hamiltonian, the EMT including the Hamiltonian density should be derivable from the Hamiltonian. In non-relativistic quantum many-body theory, the Hamiltonian density can be obtained by localizing the Hamiltonian operator, viz. inserting $\delta^{3}(r-r_{i})$ for the $i$-th constituent. For example, the charge density can be obtained by localizing the charge,
@@
In relativistic quantum theory, particles cannot be consistently localized [^2] For mildly relativistic systems, Newton-Wigner operator can be used to localize particles [168]. The application of this operator to QCD can be found in Ref. [169].. Fortunately, there is one way to circumvent this problem using light-front dynamics. In light-front dynamics, particles can be localized on the transverse plane using the Dirac-$\delta$ prescription $\delta^{2}(r_{\perp}-r_{i\perp})$, which is sufficient to extract the transverse densities. For example, the transverse charge density can be obtained by localizing the charge on the transverse plane,
@@
One can show that the charge density obtained in this way reproduces the Drell-Yan-West formula. Similarly, the matter density can also be obtained using particle localization on the light front.
@@
$M^{2}_{\text{int}}(q^{2}_{\perp})=\frac{1}{2}\sum_{s,\bar{s}}\int\frac{{\mathrm{d}}x}{4\pi x(1-x)}\int{\mathrm{d}}^{2}r_{\perp}\widetilde{\psi}^{*}_{s\bar{s}}(x,\vec{r}_{\perp})\\ \times\Big{[}e^{i\vec{q}_{\perp}\cdot\vec{r}_{1\perp}}+e^{i\vec{q}_{\perp}\cdot\vec{r}_{2\perp}}\Big{]}v(x,\vec{r}_{\perp},-i\nabla_{\perp})\widetilde{\psi}_{s\bar{s}}(x,\vec{r}_{\perp}),$ (81)
@@
$v(x,\vec{r}_{\perp},-i\nabla_{\perp})=M^{2}-\frac{-\nabla_{\perp}^{2}+m_{q}^{2}}{x}-\frac{-\nabla_{\perp}^{2}+m^{2}_{\bar{q}}}{1-x}.$ (82)
@@
Therefore, the von Laue condition is satisfied.
@@ V Numerical results
$D$
$r_{A}$ (fm)
$r_{E}$ (fm)
$r_{M^{2}}$ (fm)
$r_{\theta}$ (fm)
${2\langle T\rangle}/{\langle V\rangle}$
$\eta_{c}$
$-5.05(8)$
$0.18(1)$
$0.26(3)$
0.33(3)
0.39(1)
$-1.7(6)$
$\eta_{c}^{\prime}$
$-7.1(3)$
$0.35(2)$
$0.40(4)$
0.452(1)
0.50(2)
$-14(16)$
$\chi_{c0}$
$-6.95(2)$
$0.24(2)$
$0.29(2)$
0.39(1)
0.44(1)
$-3.0(1.1)$
@@
From these form factors, we can extract physical densities. One of the advantages of the BLFQ approach is that the coordinate-space density can be obtained in closed form. This avoids the problem of numeric fitting or numerical Fourier transform, which introduces large and uncontrollable numerical errors. The extracted pressure is shown in Fig. 4 for $\eta_{c}$, $\eta^{\prime}_{c}$ [i.e. $\eta_{c}(2S)$] and $\chi_{c0}$. The total pressure vanishes as a result of the von Laue condition. For all three states, the pressure is positive (repulsive) at the center of the meson and negative (attractive) near the edge, satisfying the stability conjecture [56]. The pressure profile of the radial excited charmonium $\eta^{\prime}_{c}$ shows one additional node. Note that the uncertainty of the pressure at small $r_{\perp}$ is large. This is because our basis has a finite UV coverage. The effective UV cutoff for $N_{\text{max}}=8$ is $\Lambda_{\text{UV}}=\kappa\sqrt{N_{\text{max}}}\approx 2.8\,\mathrm{GeV}$, corresponding to a spatial resolution of 0.07 fm. Smaller than this distance, the pressure starts to show large uncertainty, i.e. stronger basis dependence.
@@
We have mentioned the difference between the proper energy distribution $\mathcal{E}(r_{\perp})$ defined from the instant form and the invariant mass squared distribution $\mathcal{M}^{2}(r_{\perp})$ obtained on the light front. The former is normalized to the total energy of a hadron at rest, i.e. its rest mass $M$, whereas the latter is normalized to the invariant mass squared $M^{2}$ of the hadron. Fig. 5 compares the normalized energy distribution of $\eta_{c}$ and its normalized invariant mass squared distribution. The energy distribution is positive. However, the $M^{2}$ distribution becomes negative at short distance $r_{\perp}$. This tachyonic core within charmonium reveals a tantalizing novel structure of the QCD bound state. Similar quantities can be extracted and compared for $\eta^{\prime}_{c}$ and $\chi_{c0}$, as shown in Fig. 6. Within basis uncertainty, the energy distributions $\mathcal{E}(r_{\perp})$ are consistent with the weak energy condition that the energy densities are positive.
Figure 5: (Top) Comparison of the normalized energy distribution and the normalized distribution of the invariant mass squared within $\eta_{c}$. Note that the invariant mass squared density is negative at the center of $\eta_{c}$, shown in the bottom panel. This tachyonic core suggests a tantalizing picture of the strong force. The uncertainty bands are determined as described in the caption to Fig. 2.
@@
We have so far introduced several densities: the matter density $\mathcal{A}(r_{\perp})$, energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ and scalar density $\theta(r_{\perp})$. We have argued that a negative $D$ suggests a hierarchy of the radii associated with these densities:
@@
Figure 7 compares these densities for $\eta_{c}$ after the normalization, i.e. all densities are normalized to unity. Both the invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ and the scalar density $\theta(r_{\perp})$ become negative at the center of the meson. The emergence of a tachyonic core is consistent with a relatively large radius.
@@
Figure 8: (Top) Decomposition of the invariant mass squared density of $\eta_{c}$ into the free part and the interacting part. (Middle) Comparison of the free invariant mass squared densities for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$. (Bottom) Comparison of the interacting invariant mass squared densities for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$. The uncertainty bands are determined as described in the caption to Fig. 2.
As we mentioned, in the light-front Hamiltonian formalism, we are also able to access the light-front kinetic energy and light-front potential energy separately. The distributions of the free invariant mass squared as well as the interacting invariant mass squared for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$ are shown in Fig. 8. Note that $\mathcal{M}_{0}^{2}(r_{\perp})$ does not remain positive for all distance for $\eta_{c}^{\prime}$ and $\chi_{c0}$. This is caused by the recoil term $-{q^{2}_{\perp}}/{4x}$ in (67). The interaction density $\mathcal{M}_{\text{int}}^{2}(r_{\perp})$ of $\eta_{c}^{\prime}$ becomes positive at some large distance, consistent with the positive confining potential $r^{2}_{\perp}$ dominant at large parton separation.
@@
This index $n$ characterizes the shape of the inter-particle force under spatial dilation. In non-relativistic limit, our effective interaction is a harmonic oscillator $r^{2}$ at large distance plus a Coulomb $r^{-1}$ at short distance. One would expect a virial index close to $-1$ for deep bound state, e.g. $\eta_{c}$, and $+1$ for bound state with large size [170]. The obtained indices are collected Table 1. Within the sizable basis uncertainty, these numbers are in agreement with the prediction from the non-relativistic virial theorem. Fig. 9 shows the decomposition of the invariant mass squared of charmonia $\eta_{c}$, $\eta^{\prime}_{c}$ and $\chi_{c0}$ into quark mass contribution, kinetic energy contribution and the potential energy contribution.
Figure 9: Decomposition of charmonium invariant mass squared in terms of the total quark mass squared $4m_{c}^{2}$, kinetic mass squared $\langle T\rangle$ and potential mass squared $\langle V\rangle$. See texts.

### v2

@@ IV.2 Gravitational form factor $D$: kinetic part
The GFF $D(q^{2}_{\perp})$ and the light-front energy form factor $M^{2}(q^{2}_{\perp})$ can be extracted from either the light-front energy density $T^{+-}$, or the transverse stress $T^{11}+T^{22}$ and $T^{12}$. In practical model calculations, such as this work, truncations and approximations are introduced which result in the loss of some Poincaré symmetries. One then deals with a system with a reduced number of symmetries, and GFFs extracted from different current components are no longer equivalent. One possible way to extract the form factors is to parametrize the HMEs in terms of the reduced symmetry. Effectively, one introduces additional Lorentz tensor structures. Form factors associated with non-covariant structures are called spurious ones. See Ref. [49] and the references therein for a discussion on this approach to the EMT. From the covariant light-front dynamics analysis [49], $T^{11}+T^{22}$ and $T^{12}$ are contaminated by spurious form factors and may violate the von Laue condition [^1] We recently discovered a new symmetry that may protect $T^{12}$ from the contamination of the spurious form factor [172]. It is not clear whether this symmetry is applicable to the present system.. $T^{+-}$ is the light-front energy density and it contains the interaction. While this operator is more involved, it is constrained by energy conservation which ensures the von Laue condition even after model truncation. We therefore use $T^{+-}$ to extract GFF $D$. Our method is general enough and can be applied to other bound-state systems. The HME of $T^{+-}$ consists of two parts:
@@
In the Breit frame $P_{\perp}=0$, only the light-front energy form factor $M^{2}(q^{2}_{\perp})$ survives. Form factor $M^{2}(q^{2}_{\perp})$ is related to GFFs $A$ and $D$ as given in (28).
In Ref. [49], we propose to split $T^{+-}$ into two parts $T^{+-}=T^{+-}_{0}+T^{+-}_{\text{int}}$: a free part $T^{+-}_{0}$ and an interaction part $T^{+-}_{\text{int}}$. We adopt the light-cone gauge $A^{+}=0$ and in this gauge, the kinetic part $T^{+-}_{\text{0}}$ is simple. The corresponding invariant mass squared form factor is also split into two pieces, $M^{2}(q^{2}_{\perp})=M^{2}_{0}(q^{2}_{\perp})+M^{2}_{\text{int}}(q^{2}_{\perp})$. The free energy density $T^{+-}_{0}$ is diagonal in Fock space, and the corresponding invariant mass squared form factor admits an exact LFWF representation,
@@
The integrand is the $n$-body kinetic energy with an extra term $-\frac{1}{4}q^{2}_{\perp}$. This term accounts for the recoil effect, which has the same origin as the $D$-term as one can see from the free boson case [2] (cf. [49]).
@@
$M_{0}^{2}(q^{2}_{\perp})=\sum_{n}\sum_{\{s_{i}\}}\int\big{[}{\mathrm{d}}x_{i}{\mathrm{d}}^{2}r_{i\perp}\big{]}_{n}\widetilde{\psi}_{n}^{*}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\\ \times\sum_{j}e^{i\vec{r}_{j\perp}\cdot\vec{q}_{\perp}}\frac{-\vec{\nabla}_{j\perp}^{2}+m_{j}^{2}-\frac{1}{4}q_{\perp}^{2}}{x_{j}}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\,.$ (66)
Its Fourier transform gives the OBD of the free light-front energy,
$\mathcal{M}_{0}^{2}(r_{\perp})=\sum_{n}\sum_{\{s_{i}\}}\int\big{[}{\mathrm{d}}x_{i}{\mathrm{d}}^{2}r_{i\perp}\big{]}_{n}\sum_{j}\delta^{2}(\vec{r}_{j\perp}-\vec{r}_{\perp})\\ \times\bigg{\{}\widetilde{\psi}_{n}^{*}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\frac{-\vec{\nabla}_{j\perp}^{2}+m_{j}^{2}}{x_{j}}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\\ +\frac{1}{4x_{j}}\nabla^{2}_{j\perp}\big{|}\widetilde{\psi}_{n}(\{x_{i},\vec{r}_{i\perp},s_{i}\})\big{|}^{2}\bigg{\}}\,.$ (67)
@@ IV.3 Gravitational form factor $D$: interaction part
The interaction part of the EMT in coordinate space can be formally represented in Fock space using LFWFs. However, the expression is not particularly useful in quantum field theories, since it involves particle number changing. Furthermore, the interaction $v_{nm}$ also has to be renormalized. The effective interactions, such as the one-gluon-exchange we employed, are non-local in general. It may seem that a local interaction density has to be obtained from an underlying field theory. On the other hand, in non-relativistic quantum many-body theory, the Hamiltonian density can also be obtained by localizing the Hamiltonian operator, viz. inserting $\delta^{3}(r-r_{i})$ for the $i$-th constituent. For example, the charge density can be obtained by localizing the charge,
@@
The above method can be generalized to the relativistic quantum many-body systems. In relativistic quantum theory, particles cannot be consistently localized [^2] For mildly relativistic systems ($v/c\sim 0.1$), Newton-Wigner operator can be used to localize particles [175]. The application of this operator to QCD can be found in Ref. [176].. Fortunately, there is one way to circumvent this problem using light-front dynamics. In light-front dynamics, particles can be localized on the transverse plane using the Dirac-$\delta$ prescription $\delta^{2}(r_{\perp}-r_{i\perp})$, which is sufficient to extract the transverse densities. For example, the transverse charge density can be obtained by localizing the charge on the transverse plane,
@@
One can show that the charge density obtained in this way reproduces the Drell-Yan-West formula [173, 174]. Similarly, the matter density can also be obtained using particle localization on the light front.
@@
$M^{2}_{\text{int}}(q^{2}_{\perp})=\frac{1}{2}\sum_{s,\bar{s}}\int\frac{{\mathrm{d}}x}{4\pi x(1-x)}\int{\mathrm{d}}^{2}r_{\perp}\widetilde{\psi}^{*}_{s\bar{s}}(x,\vec{r}_{\perp})\\ \times\Big{[}e^{i\vec{q}_{\perp}\cdot\vec{r}_{1\perp}}+e^{i\vec{q}_{\perp}\cdot\vec{r}_{2\perp}}\Big{]}v(x,\vec{r}_{\perp},-i\vec{\nabla}_{\perp})\widetilde{\psi}_{s\bar{s}}(x,\vec{r}_{\perp}),$ (80)
@@
$v(x,\vec{r}_{\perp},-i\vec{\nabla}_{\perp})=M^{2}-\frac{-\nabla_{\perp}^{2}+m_{q}^{2}}{x}-\frac{-\nabla_{\perp}^{2}+m^{2}_{\bar{q}}}{1-x}.$ (81)
@@
Therefore, the forward limit constraint is satisfied.
@@ V Numerical results
$D$ $r_{A}$ (fm) $r_{E}$ (fm) $r_{M^{2}}$ (fm) $r_{\theta}$ (fm) ${2\langle T\rangle}/{\langle V\rangle}$
$\eta_{c}$ $-5.05(8)$ $0.18(1)$ $0.26(3)$ 0.33(3) 0.39(1) $-1.7(6)$
$\eta_{c}^{\prime}$ $-7.1(3)$ $0.35(2)$ $0.40(4)$ 0.452(1) 0.50(2) $-14(16)$
$\chi_{c0}$ $-6.95(2)$ $0.24(2)$ $0.29(2)$ 0.39(1) 0.44(1) $-3.0(1.1)$
@@
From these form factors, we can extract physical densities. One of the advantages of the BLFQ approach is that the coordinate-space density can be obtained in closed form. This avoids the problem of numeric fitting or numerical Fourier transform, which introduces large and uncontrollable numerical errors. The extracted pressure is shown in Fig. 4 for $\eta_{c}$, $\eta^{\prime}_{c}$ [i.e. $\eta_{c}(2S)$] and $\chi_{c0}$. The total pressure vanishes as a result of the von Laue condition. For all three states, the pressure is positive (repulsive) at the center of the meson and negative (attractive) near the edge, satisfying the stability conjecture [167, 168]. The pressure profile of the radial excited charmonium $\eta^{\prime}_{c}$ shows two additional nodes [23]. Note that the uncertainty of the pressure at small $r_{\perp}$ is large. This is because our basis has a finite UV coverage. The effective UV cutoff for $N_{\text{max}}=8$ is $\Lambda_{\text{UV}}=\kappa\sqrt{N_{\text{max}}}\approx 2.8\,\mathrm{GeV}$, corresponding to a spatial resolution of 0.07 fm. Smaller than this distance, the pressure starts to show large uncertainty, i.e. stronger basis dependence.
@@
We have mentioned the difference between the proper energy distribution $\mathcal{E}(r_{\perp})$ defined from the instant form and the invariant mass squared distribution $\mathcal{M}^{2}(r_{\perp})$ obtained on the light front. The former is normalized to the total energy of a hadron at rest, i.e. its rest mass $M$, whereas the latter is normalized to the invariant mass squared $M^{2}$ of the hadron. Fig. 5 compares the normalized energy distribution of $\eta_{c}$ and its normalized invariant mass squared distribution. The energy distribution is positive, consistent with the weak energy condition. However, the invariant mass squared distribution becomes negative (i.e. tachyonic) at short distance $r_{\perp}$, although at the short distance, the results suffer from slow basis convergence as indicated by the wide bands. Similar quantities can be extracted and compared for $\eta^{\prime}_{c}$ and $\chi_{c0}$, as shown in Fig. 6. Within basis uncertainty, the energy distributions $\mathcal{E}(r_{\perp})$ are consistent with the weak energy condition that the energy densities are positive.
Figure 5: Comparison of the normalized energy distribution and the normalized distribution of the invariant mass squared within $\eta_{c}$. The uncertainty bands are determined as described in the caption to Fig. 2.
@@
We have so far introduced several densities: the matter density $\mathcal{A}(r_{\perp})$, energy density $\mathcal{E}(r_{\perp})$, invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ and (trace) scalar density $\theta(r_{\perp})$. Both $\mathcal{E}(r_{\perp})$ and $\mathcal{A}(r_{\perp})$ are strictly positive as required by the weak and null energy conditions. The energy conditions impose constraints on $D$ (46), which are satisfied by a negative $D$. We have argued that a negative $D$ suggests a hierarchy of the radii associated with these densities:
@@
Figure 7 compares these densities for $\eta_{c}$ after the normalization, i.e. all densities are normalized to unity. Both the invariant mass squared density $\mathcal{M}^{2}(r_{\perp})$ and the scalar density $\theta(r_{\perp})$ become negative at the center of the meson.
@@
Figure 8: (Top) Decomposition of the internal light-front energy density, i.e. invariant mass squared density, of $\eta_{c}$ into the free part and the interacting part. (Middle) Comparison of the free invariant mass squared densities for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$. (Bottom) Comparison of the interacting invariant mass squared densities for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$. The uncertainty bands are determined as described in the caption to Fig. 2.
As we mentioned, in the light-front Hamiltonian formalism, we are also able to access the light-front kinetic energy and light-front potential energy separately. The distributions of the free internal light-front energy (i.e. free invariant mass squared) as well as the interacting light-front energy for $\eta_{c}$, $\eta_{c}^{\prime}$ and $\chi_{c0}$ are shown in Fig. 8. Note that $\mathcal{M}_{0}^{2}(r_{\perp})$ does not remain positive for all distance for $\eta_{c}^{\prime}$ and $\chi_{c0}$. This is caused by the recoil term $-{q^{2}_{\perp}}/{4x}$ in (66). The interaction density $\mathcal{M}_{\text{int}}^{2}(r_{\perp})$ of $\eta_{c}^{\prime}$ becomes positive at some large distance, consistent with the positive confining potential $r^{2}_{\perp}$ dominant at large parton separation.
@@
This index $n$ characterizes the shape of the inter-particle force under spatial dilation. In non-relativistic limit, our effective interaction is a harmonic oscillator $r^{2}$ at large distance plus a Coulomb $r^{-1}$ at short distance. One would expect a virial index close to $-1$ for deep bound state, e.g. $\eta_{c}$, and $+1$ for bound state with large size [178]. The obtained indices are collected Table 1. Within the sizable basis uncertainty, these numbers are in agreement with the prediction from the non-relativistic virial theorem. Fig. 9 shows the decomposition of the invariant mass squared $M^{2}$ of charmonia $\eta_{c}$, $\eta^{\prime}_{c}$ and $\chi_{c0}$ into quark mass contribution, kinetic energy contribution and the potential energy contribution.
Figure 9: Decomposition of charmonium invariant mass squared $M^{2}$ in terms of the total quark mass squared $4m_{c}^{2}$, kinetic light-front energy $\langle T\rangle$ and potential light-front energy $\langle V\rangle$. See texts.